# references.bib

### v1

@@ 5.2 Boundary terms of Lindbladian form
We address this question by constructing the double row transfer matrix, as given in (2.11) and, by taking the higher order derivatives, we calculate the analytical expression of the terms at the boundary. By taking the higher order derivative of the double row transfer matrix, one can prove that the first non-trivial conserved charge is
$\mathbb{Q}_{2}=\sum q^{(2)}_{i,i+1}+q^{(2),L}_{L}+q^{(2),R}_{1}.$ (5.9)
@@
$q^{(2),L/R}_{L/1}=\sum_{i}\Big{[}\ell^{(i)}_{L/1}\otimes\ell^{(i)*}_{L/1}-\frac{1}{2}\big{(}\ell^{(i)\dagger}_{L/1}\ell^{(i)}_{L/1}\big{)}\otimes\mathbb{I}-\frac{1}{2}\mathbb{I}\otimes\big{(}\ell^{(i)t}_{L/1}\ell^{(i)*}_{L/1}\big{)}\Big{]},$ (5.10)
@@
First, we have to obtain the analytical expression of $q^{(2),L/R}_{L/1}$.
@@ 5.2.1 Analytical expression of the boundary terms
For the class of models we are considering, the first and second derivative of the transfer matrix do not generate the dynamics. In fact, the first derivative vanishes and the second derivative is proportional to the identity operator. The first non-trivial charge is the derivative of order 3 of the transfer matrix. For completeness, since we were not able to find this expressions in the literature, we give the analytical expression of the first two charges in the appendix C. We derived these expressions, as well as the one for the third derivatives, using Mathematica software and the non-commutative product. The precise expression for the third derivative in a general model is quite lengthy, so it is not included in this paper. However, upon request, we are happy to share the Mathematica notebook with anyone interested. For the model we are considering, many terms are zero and the final expression is very compact.
We obtained
$\partial_{u}^{3}t(u)|_{u\to 0}=\big{(}\text{tr}(\partial^{2}_{u}K^{L}(0))\big{)}q^{(2)}_{Bulk}+\frac{1}{2}\big{(}\text{tr}(\partial^{2}_{u}K^{L}(0))\big{)}\partial_{u}K^{R}_{1}(0)+2\,\text{tr}_{0}(\partial_{u}K^{L}_{0}(0)h_{0L}^{2}).$ (5.11)
@@
$q^{(2)}=q^{(2)}_{Bulk}+q^{(2),L}_{L}+q^{(2),R}_{1},\,\,\,\,\,\text{where}\,\,\,\,\,q^{(2),L}_{L}=2\,\frac{\text{tr}_{0}(\partial_{u}K^{L}_{0}(0)h_{0L}^{2})}{\text{tr}(\partial^{2}_{u}K^{L}(0))}\,\,\text{and}\,\,q^{(2),R}_{1}=\frac{1}{2}\partial_{u}K^{R}_{1}(0).$ (5.12)
@@ 5.2.2 Attempt to write the solution as Lindbladian
Having found the analytical expression, we can now check whether the boundary terms $q^{(2),L}_{L}$ and $q^{(2),R}_{1}$ can be brought into a Lindbladian form (5.10).
@@
$\ell^{(i)}=\left(\begin{array}[]{cc}\lambda_{i,1}&\lambda_{i,2}\\ \lambda_{i,3}&\lambda_{i,4}\\ \end{array}\right)$ (5.13)
and find a solutions for the $\lambda_{i,j}$ such that (5.10) is satisfied. However, this approach will be computationally very hard since the sum runs over the square of dimension of the local Hilbert space.
@@
First, we calculated the boundary terms corresponding to the XX spin chain given in section 4.2.1 and to the non factorized solution (4.23) and we obtain
$q_{1}^{(2),R}\propto\left(\begin{array}[]{cccc}\kappa_{1,1}&0&0&\kappa_{1,4}\\ 0&0&\kappa_{2,3}&0\\ 0&\kappa_{3,2}&\kappa_{3,3}&0\\ \kappa_{4,1}&0&0&\kappa_{3,3}-\kappa_{1,1}\\ \end{array}\right)$ (5.14)
@@
$q_{L}^{(2),L}\propto\left(\begin{array}[]{cccc}2\tilde{\kappa}_{1,1}-\tilde{\kappa}_{3,3}&0&0&2\tilde{\kappa}_{1,4}\\ 0&-\tilde{\kappa}_{3,3}&2\tilde{\kappa}_{2,3}&0\\ 0&2\tilde{\kappa}_{3,2}&\tilde{\kappa}_{3,3}&0\\ 2\tilde{\kappa}_{4,1}&0&0&\tilde{\kappa}_{3,3}-2\tilde{\kappa}_{1,1}\\ \end{array}\right).$ (5.15)
@@
$k_{1}+k_{4}=k_{2}+k_{3}.$ (5.16)
First, we construct the matrix form of the operators $q^{(2),L/R}_{L/1}$ in (5.10), by using the definition (5.13) for the $\ell^{(i)}$. The condition (5.16) for the operator $\ell^{(i)}$ in (5.10) implies
$\sum_{i}|\lambda_{i,1}-\lambda_{i,4}|^{2}=0,$ (5.17)
@@
$\lambda_{i,1}=\lambda_{i,4},\,\,\,\,\,\forall\,i,$ (5.18)
@@
This proves that the condition (5.16) could be satisfied only if $\ell^{(i)}$ has the structure
$\ell^{(i)}=\left(\begin{array}[]{cc}0&\lambda_{i,2}\\ \lambda_{i,3}&0\\ \end{array}\right)=\lambda_{i,2}\sigma^{+}+\lambda_{i,3}\sigma^{-}.$ (5.21)
At this point, to match the expression (5.10) is much easier and we obtain that, if we set
$\kappa_{3,2}=\kappa_{2,3},$ $\kappa_{1,1}=-\kappa_{4,1},$ $\kappa_{3,3}=0,$ $\kappa_{1,4}=-\kappa_{4,1}$ (5.22)
@@
$\tilde{\kappa}_{1,4}=\tilde{\kappa}_{4,1},$ $\tilde{\kappa}_{3,2}=\tilde{\kappa}_{2,3},$ $\tilde{\kappa}_{3,3}=0,$ $\tilde{\kappa}_{1,1}=0,$ (5.23)
we can rescale the remaining free parameters such that the right and left boundaries are described by
$\ell_{1}^{(1)}=\alpha\sigma^{+},$ $\ell_{1}^{(2)}=\beta\sigma^{-},$ $\ell_{1}^{(3)}=\gamma_{1}\sigma^{+}+\gamma_{2}\sigma^{-}$ (5.24)
and similarly
$\ell_{L}^{(1)}=\tilde{\alpha}\sigma^{+},$ $\ell_{L}^{(2)}=\tilde{\beta}\sigma^{-},$ $\ell_{L}^{(3)}=\tilde{\gamma}_{1}\sigma^{+}+\tilde{\gamma}_{2}\sigma^{-},$ (5.25)
where $\alpha$, $\beta$, $\gamma_{1}$ and $\gamma_{2}$ (and their tilde versions) are free constants.
We conclude that, the XX spin chain in the bulk, with boundary terms described by the jump operators (5.24) and (5.25) is Yang-Baxter integrable.
5.2.3 XX spin chain with magnetic field
We consider the XX spin chain with magnetic field given in section 4.2.2, we calculate the boundary term for the non factorized solution (4.36) and we obtain
$q_{1}^{(2),R}\propto\left(\begin{array}[]{cccc}\kappa_{3,3}-\kappa_{4,4}&0&0&\kappa_{1,4}\\ 0&0&0&0\\ 0&0&\kappa_{3,3}&0\\ \kappa_{4,1}&0&0&\kappa_{4,4}\\ \end{array}\right)$ (5.30)
@@ 5.2.3 XX spin chain with magnetic field
$q_{L}^{(2),L}\propto\left(\begin{array}[]{cccc}2\tilde{\kappa}_{4,4}\csc^{2}\psi-\tilde{\kappa}_{3,3}&0&0&-2\tilde{\kappa}_{4,1}\csc^{2}\psi\\ 0&\tilde{\kappa}_{3,3}&0&0\\ 0&0&\tilde{\kappa}_{3,3}\left(2\csc^{2}\psi-3\right)&0\\ -2\tilde{\kappa}_{1,4}\csc^{2}\psi&0&0&\tilde{\kappa}_{3,3}\left(2\csc^{2}\psi-1\right)-2\tilde{\kappa}_{4,4}\csc^{2}\psi\\ \end{array}\right).$ (5.35)
$\kappa_{3,3}=0,$ $\kappa_{4,4}=-\frac{1}{2}\left(\kappa_{1,4}-\kappa_{4,1}\right)$ (5.36)
and
$\tilde{\kappa}_{3,3}=0,$ $\tilde{\kappa}_{4,4}=\frac{1}{2}\left(\tilde{\kappa}_{1,4}-\tilde{\kappa}_{4,1}\right).$ (5.37)
@@
$\ell_{1}^{(1)}=\alpha\,\sigma^{+},$ $\ell_{1}^{(2)}=\beta\,\sigma^{-}$ (5.38)
@@
$\ell_{L}^{(1)}=\tilde{\alpha}\,\sigma^{+},$ $\ell_{L}^{(2)}=\tilde{\beta}\,\sigma^{-},$ (5.39)
@@
We conclude that, the XX spin chain with magnetic field in the bulk, with boundary terms described by the jump operators (5.38) and (5.39) is Yang-Baxter integrable.
@@
$q^{(2),R}=\partial_{u}K^{(2),R}(u)|_{u=0}=\left(\begin{array}[]{cccc}\kappa_{3,3}-\kappa_{4,4}&0&0&\kappa_{1,4}\\ 0&0&0&0\\ 0&0&\kappa_{3,3}&0\\ \kappa_{4,1}&0&0&\kappa_{4,4}\\ \end{array}\right)$ (5.45)
@@
$q^{(2),L}=\left(\begin{array}[]{cccc}\tilde{\kappa}_{3,3}-\tilde{\kappa}_{4,4}&0&0&\tilde{\kappa}_{1,4}\\ 0&0&0&0\\ 0&0&\tilde{\kappa}_{3,3}&0\\ \tilde{\kappa}_{4,1}&0&0&\tilde{\kappa}_{4,4}\\ \end{array}\right).$ (5.50)
The expression for the boundaries coincide with the ones in (5.30), so in order to bring the solution to a Lindbladian form, we have to require the condition (5.36) for the coefficients and the expressions for the $\ell$ operators are (5.38) and (5.39).
It is worth to mention that for this model, an interesting phenomenon happens. The first non-trivial conserved charge is generated by taking the first derivative of the transfer matrix, however, the analytical expression does not coincide with (26) given by Sklyanin in [Sklyanin:1987bi]. The reason is that, in Sklyanin’s derivation, he assumes that taking the limit of the product of matrices as the spectral parameter approaches zero is equivalent to taking the limit of each term in the product individually. However, there can be cases, such as the one we are considering, where this assumption does not hold. In our case, one factor diverges in the limit. Taking the product first regularizes the result. We comment further on this statement in Appendix C.
@@ 6 Conclusion
In this work, we constructed quantum circuits on a qubit ladder where the elementary quantum gate in the bulk is a tensor product of a pair of two qubit gates. We investigate whether the quantum gates at the left and right boundaries may exhibit a non-factorized expression while preserving integrability. We obtain that non-factorized solutions exist if the bulk model is the free fermion XX spin chain. In contrast, for the other models analyzed, namely the XXX and XXZ spin chains, all solutions are factorized. For the free fermion model, we examine the continuous time dynamics and find that the boundary terms can be expressed as a Lindbladian evolution for a specific choice of free parameters. This means that the operators at the left and right boundaries of the spin chain act as sources for particle injection or removal. This result implies that the Yang-Baxter integrability property of a model, both in the bulk and at the boundaries, imposes stricter conditions than those required to find the analytical expression of the NESS of the model.
This result is somehow reminiscent to a similar phenomenon in quantum field theory (QFT) with extended defect lines or with an impurity, [delfino1994scattering, castro2002absence]. There it was observed that only integrable QFTs with diagonal $S$-matrix in the bulk and factorizable scattering exhibiting both transmission and reflection are those associated to generalized-free theories.
Many interesting open questions need to be addressed. One may introduce the coupling between the two free fermionic (XX) chains either in the form of a Hubbard interaction in the unitary context or dephasing in the open systems framework, and again ask the question about factorizability or Lindbladian form of the boundary terms. Furthermore, we observed that the circuit of Fig. 10, corresponds to a finite free fermionic model with open boundaries and an interacting impurity in the middle. For a particular choice of the parameters of the impurity, one has $U(1)$ symmetry which allows to study impurity related transport properties of the model. Another interesting direction is to study factorizability of $K$-matrices for higher local Hilbert space dimension.
Acknowledgments
We would like to thank Z. Bajnok, M. de Leeuw, E. Ilievski, Y. Kasim, A. Klümper, R. Nepomechie, V. Popkov, A. L. Retore, A. Torrielli, L. Zadnik, M. Žnidarič for useful discussions. We also thank A. L. Retore for useful comments on the manuscript. C.P. would like to thank P. Vieira for the kind hospitality at ICTP SAIFT in São Paulo during the last stage of this work.

### v2

@@ 5.2 Boundary terms of Lindbladian form
We address this question by constructing the double row transfer matrix, as given in (12) and, by taking the higher order derivatives and evaluate it at $u=0$, we calculate the analytical expression of the terms at the boundary. By taking the higher order derivative [^17] We remark that for the models we are considering (except for the XX spin chain with magnetic field and $h=1$), this charge corresponds to the third derivatives of $t(u)$. of the double row transfer matrix, one can prove that the first non-trivial conserved charge is
$\mathbb{Q}_{2}=\sum q^{(2)}_{i,i+1}+q^{(2),L}_{L}+q^{(2),R}_{1}.$ (106)
@@
$q^{(2),L/R}_{L/1}=\sum_{i}\Big{[}\ell^{(i)}_{L/1}\otimes\ell^{(i)*}_{L/1}-\frac{1}{2}\big{(}\ell^{(i)\dagger}_{L/1}\ell^{(i)}_{L/1}\big{)}\otimes\mathbb{I}-\frac{1}{2}\mathbb{I}\otimes\big{(}\ell^{(i)t}_{L/1}\ell^{(i)*}_{L/1}\big{)}\Big{]},$ (107)
@@
First, we have to obtain the analytical expression of the boundary terms $q^{(2),L/R}_{L/1}$ in terms of the quantities $K^{L/R}$ and the Hamiltonian $h$, or, more generally, the $R$-matrix.
@@ 5.2.1 Analytical expression of the boundary terms
For the class of models we are considering, except model of section 4.2.2 with $h=1$, the first and second derivative of the transfer matrix do not generate the dynamics. In fact, the first derivative vanishes and the second derivative is proportional to the identity operator. The first non-trivial charge is the derivative of order 3 of the transfer matrix. For completeness, since we were not able to find this expressions in the literature, we give the analytical expression of the first two charges in the appendix C. We derived these expressions, as well as the one for the third derivatives, using Mathematica software and the non-commutative product. The precise expression for the third derivative in a general model is quite lengthy, so it is not included in this paper. However, upon request, we are happy to share the Mathematica notebook with anyone interested. For the model we are considering, many terms are zero and the final expression is very compact.
We obtained, by ignoring the term proportional to the identity matrix
$\partial_{u}^{3}t(u)|_{u\to 0}=\alpha\,q^{(2)}_{Bulk}+\frac{\alpha}{2}\,\partial_{u}K^{R}_{1}(0)+\,\text{tr}_{0}\Big{(}2\,\partial_{u}K^{L}_{0}(0)\,h_{0L}^{2}+\partial_{u}^{2}K^{L}_{0}(0)\,h_{0L}+K^{L}_{0}(0)\,\partial_{u}^{3}\check{R}_{0L}(0)\Big{)},$ (108)
where
$\alpha=\text{tr}(\partial^{2}_{u}K^{L}(0))+4\,\text{tr}_{0}(K_{0}^{L}(0)\,h_{0L}^{2})+4\,\text{tr}_{0}(\partial_{u}K_{0}^{L}(0)\,h_{0L}).$ (109)
@@
$q^{(2)}=q^{(2)}_{Bulk}+q^{(2),L}_{L}+q^{(2),R}_{1},$ (110)
where
$q^{(2),L}_{L}=\frac{1}{\alpha}\text{tr}_{0}\Big{(}2\,\partial_{u}K^{L}_{0}(0)\,h_{0L}^{2}+\partial_{u}^{2}K^{L}_{0}(0)\,h_{0L}+K^{L}_{0}(0)\,\partial_{u}^{3}\check{R}_{0L}(0)\Big{)},$ (111)
$q^{(2),R}_{1}=\frac{1}{2}\partial_{u}K^{R}_{1}(0).$ (112)
@@ 5.2.2 Attempt to write the solution as Lindbladian
Having found the analytical expression, we can now check whether the boundary terms $q^{(2),L}_{L}$ and $q^{(2),R}_{1}$ can be brought into a Lindbladian form (107).
@@
$\ell^{(i)}=\left(\begin{array}[]{cc}\lambda_{i,1}&\lambda_{i,2}\\ \lambda_{i,3}&\lambda_{i,4}\\ \end{array}\right)$ (113)
and find a solutions for the $\lambda_{i,j}$ such that the decomposition (107) matches (111)-(112) separately for the right and the left boundaries. However, this approach will be computationally very hard since the sum over $i$ in (107) runs over the square of dimension of the local Hilbert space.
@@
First, we calculated the boundary terms corresponding to the XX spin chain given in section 4.2.1 and to the non factorized solution (62) and we obtain
$q_{1}^{(2),R}\propto\left(\begin{array}[]{cccc}\kappa_{1,1}&0&0&\kappa_{1,4}\\ 0&0&\kappa_{2,3}&0\\ 0&\kappa_{3,2}&\kappa_{3,3}&0\\ \kappa_{4,1}&0&0&\kappa_{3,3}-\kappa_{1,1}\\ \end{array}\right)$ (114)
@@
$q_{L}^{(2),L}\propto\left(\begin{array}[]{cccc}2\tilde{\kappa}_{1,1}-\tilde{\kappa}_{3,3}&0&0&2\tilde{\kappa}_{1,4}\\ 0&-\tilde{\kappa}_{3,3}&2\tilde{\kappa}_{2,3}&0\\ 0&2\tilde{\kappa}_{3,2}&\tilde{\kappa}_{3,3}&0\\ 2\tilde{\kappa}_{4,1}&0&0&\tilde{\kappa}_{3,3}-2\tilde{\kappa}_{1,1}\\ \end{array}\right).$ (115)
@@
$k_{1}+k_{4}=k_{2}+k_{3}.$ (116)
First, we construct the matrix form of the operators $q^{(2),L/R}_{L/1}$ in (107), by using the definition (113) for the $\ell^{(i)}$. The condition (116) for the operator $\ell^{(i)}$ in (107) implies
$\sum_{i}|\lambda_{i,1}-\lambda_{i,4}|^{2}=0,$ (117)
@@
$\lambda_{i,1}=\lambda_{i,4},\,\,\,\,\,\forall\,i,$ (118)
@@
This proves that the condition (116) could be satisfied only if $\ell^{(i)}$ has the structure
$\ell^{(i)}=\left(\begin{array}[]{cc}0&\lambda_{i,2}\\ \lambda_{i,3}&0\\ \end{array}\right)=\lambda_{i,2}\sigma^{+}+\lambda_{i,3}\sigma^{-}.$ (121)
At this point, to match the expression (107) is much easier and we obtain that, if we set [^18] We notice that the condition $\kappa_{3,3}=0$ is compatible with the fact that the dissipator in (107) is traceless.
$\kappa_{3,2}=\kappa_{2,3}^{*},$ $\kappa_{1,1}=\frac{1}{2}(\kappa_{1,4}-\kappa_{4,1}),$ $\kappa_{3,3}=0,$ (122)
and similar expressions for the $\tilde{\kappa}$, we can rescale the remaining free parameters such that the right and left boundaries are described by
$\ell_{1}^{(1)}=\alpha\sigma^{+},$ $\ell_{1}^{(2)}=\beta_{1}\sigma^{+}+\beta_{2}\sigma^{-}$ (123)
and similarly
$\ell_{L}^{(1)}=\tilde{\alpha}\sigma^{+},$ $\ell_{L}^{(2)}=\tilde{\beta}_{1}\sigma^{+}+\tilde{\beta}_{2}\sigma^{-},$ (124)
where $\alpha$, $\beta_{1}$ and $\beta_{2}$ (and their tilde versions) are free constants [^19] We remark that equivalently one can use the jump operators $\ell_{1/L}^{(2)}$ as given in (123) and (124) and $\ell_{1/L}^{(1)}$ to be proportional to $\sigma^{-}$..
We conclude that, the XX spin chain in the bulk, with boundary terms described by the jump operators (123) and (124) is Yang-Baxter integrable.
5.2.3 XX spin chain with magnetic field
We consider the XX spin chain with magnetic field given in section 4.2.2, we calculate the boundary term for the non factorized solution (75) and we obtain
$q_{1}^{(2),R}\propto\left(\begin{array}[]{cccc}\kappa_{3,3}-\kappa_{4,4}&0&0&\kappa_{1,4}\\ 0&0&0&0\\ 0&0&\kappa_{3,3}&0\\ \kappa_{4,1}&0&0&\kappa_{4,4}\\ \end{array}\right)$ (129)
@@ 5.2.3 XX spin chain with magnetic field
$q_{L}^{(2),L}\propto\left(\begin{array}[]{cccc}\frac{\tilde{\kappa}_{3,3}}{2}-\tilde{\kappa}_{4,4}\csc^{2}\psi&0&0&\tilde{\kappa}_{4,1}\csc^{2}\psi\\ 0&\zeta&0&0\\ 0&0&-\tilde{\kappa}_{3,3}\cot^{2}\psi-\zeta&0\\ \tilde{\kappa}_{1,4}\csc^{2}\psi&0&0&\tilde{\kappa}_{4,4}\csc^{2}\psi+\tilde{\kappa}_{3,3}\left(\frac{1}{2}-\csc^{2}\psi\right)\\ \end{array}\right),$ (134)
with
$\zeta=\tilde{\kappa}_{3,3}\left(\tilde{\kappa}_{4,4}\cot\psi-\frac{1}{2}\right)-\left(\tilde{\kappa}_{4,4}^{2}+\tilde{\kappa}_{1,4}\tilde{\kappa}_{4,1}+1\right)\cot\psi.$ (135)
$\kappa_{3,3}=0,$ $\kappa_{4,4}=-\frac{1}{2}\left(\kappa_{1,4}-\kappa_{4,1}\right)$ (136)
@@
$\tilde{\kappa}_{3,3}=\frac{\csc^{2}\psi\left(\left(\tilde{\kappa}_{4,4}^{2}+\tilde{\kappa}_{1,4}\tilde{\kappa}_{4,1}+1\right)\sin(2\psi)+\tilde{\kappa}_{1,4}-\tilde{\kappa}_{4,1}-2\tilde{\kappa}_{4,4}\right)}{2\tilde{\kappa}_{4,4}\cot\psi-2},$ (137)
$\tilde{\kappa}_{4,4}=\frac{1}{4}\left(-\csc\psi\sec\psi\sqrt{\left(\tilde{\kappa}_{1,4}+\tilde{\kappa}_{4,1}\right){}^{2}\sin^{2}(2\psi)+4}+2\tilde{\kappa}_{1,4}-2\tilde{\kappa}_{4,1}-4\cot(2\psi)\right).$ (138)
@@
$\ell_{1}^{(1)}=\alpha\,\sigma^{+},$ $\ell_{1}^{(2)}=\beta\,\sigma^{-}$ (139)
@@
$\ell_{L}^{(1)}=\tilde{\alpha}\,\sigma^{+},$ $\ell_{L}^{(2)}=\tilde{\beta}\,\sigma^{-},$ (140)
@@
We conclude that, the XX spin chain with magnetic field in the bulk, with boundary terms described by the jump operators (139) and (140) is Yang-Baxter integrable.
@@
$q^{(2),R}=\partial_{u}K^{(2),R}(u)|_{u=0}=\left(\begin{array}[]{cccc}\kappa_{3,3}-\kappa_{4,4}&0&0&\kappa_{1,4}\\ 0&0&0&0\\ 0&0&\kappa_{3,3}&0\\ \kappa_{4,1}&0&0&\kappa_{4,4}\\ \end{array}\right)$ (146)
@@
$q^{(2),L}=\left(\begin{array}[]{cccc}\tilde{\kappa}_{3,3}-\tilde{\kappa}_{4,4}&0&0&\tilde{\kappa}_{1,4}\\ 0&0&0&0\\ 0&0&\tilde{\kappa}_{3,3}&0\\ \tilde{\kappa}_{4,1}&0&0&\tilde{\kappa}_{4,4}\\ \end{array}\right).$ (151)
The expression for the boundaries coincide with the ones in (129), so in order to bring the solution to a Lindbladian form, we have to require the condition (136) for the coefficients (both $\kappa$s and $\tilde{\kappa}$s) and the expressions for the $\ell$ operators are (139) and (140).
It is worth to mention that for this model, an interesting phenomenon happens. The first non-trivial conserved charge is generated by taking the first derivative of the transfer matrix, however, the analytical expression does not coincide with (26) given by Sklyanin in [45]. The reason is that, in Sklyanin’s derivation, he assumes that taking the limit of the product of matrices as the spectral parameter approaches zero is equivalent to taking the limit of each term in the product individually. However, there can be cases, such as the one we are considering, where this assumption does not hold. In our case, one factor diverges in the limit. Taking the product first regularizes the result. We comment further on this statement in Appendix C.
@@ 6 Conclusion
In this work, we constructed quantum circuits on a qubit ladder where the elementary quantum gate in the bulk is a tensor product of a pair of two qubit gates. We investigate whether the quantum gates at the left and right boundaries may exhibit a non-factorized expression while preserving integrability. We obtain that non-factorized solutions exist only if the bulk model is the free fermion XX spin chain. In contrast, for the other models analyzed, namely the XXX and XXZ spin chains, all solutions are factorized. For the free fermion model, we examine the continuous time dynamics and find that the boundary terms can be expressed as a Lindbladian evolution for a specific choice of free parameters. This means that the operators at the left and right boundaries of the spin chain act as sources for particle injection or removal. This result implies that the Yang-Baxter integrability property of a model, both in the bulk and at the boundaries, imposes stricter conditions than those required to find the analytical expression of the NESS of the model.
The free fermionic property of the model in the bulk seems to be also a necessary condition to guarantee the Yang-Baxter integrability of the Lindbladian superoperator with dissipator terms acting on the bulk of the spin chain, [19, 46, 47].
This result is somehow reminiscent to a similar phenomenon in quantum field theory (QFT) with extended defect lines or with an impurity, [48, 49]. There it was observed that only integrable QFTs with diagonal $S$-matrix in the bulk and factorizable scattering exhibiting both transmission and reflection are those associated to generalized-free theories.
Many interesting open questions need to be addressed. One may introduce the coupling between the two free fermionic (XX) chains either in the form of a Hubbard interaction in the unitary context or dephasing in the open systems framework, and again ask the question about factorizability or Lindbladian form of the boundary terms. Furthermore, we observed that the circuit of Fig. 10 where the gate is the XX model, corresponds to a finite free fermionic model with open boundaries and an interacting impurity in the middle. For a particular choice of the parameters of the impurity, one has $U(1)$ symmetry which allows to study impurity related transport properties of the model. Another interesting direction is to study factorizability of $K$-matrices for higher local Hilbert space dimension.
Acknowledgements
We would like to thank Z. Bajnok, M. de Leeuw, E. Ilievski, Y. Kasim, A. Klümper, R. Nepomechie, V. Popkov, A. L. Retore, A. Torrielli, L. Zadnik, M. Žnidarič for useful discussions. We also thank A. L. Retore for useful comments on the manuscript. C.P. would like to thank P. Vieira for the kind hospitality at ICTP SAIFR in São Paulo during the last stage of this work.